OpenStudy (anonymous):

you determine that it is feasible to sell your specialty pizzas for $15. The cost for making the pizzas includes a fixed cost of $55 and a labor cost of $4 per pizza. •Establish an equation to determine revenue. •Establish an equation to determine total cost. •How many pizzas must be sold to break even (i.e., you experience neither a profit nor a loss)? As a business owner, it is very important to have an understanding of profit and loss. The formula for determining profit is Profit = revenue – cost (P = R – C). •Determine the profit if 500 specialty pizzas are sold. •How many spe

OpenStudy (hoblos):

let x= number of pizza R=15x C=55+4x break even---> 15x=55+4x 11x=55 x=5 so 5 pizzas must be sold to break even P = R – C = 15x - (55+4x) = 11x -55 for x=500 P = 5500 -55 = 5445

OpenStudy (hoblos):

the last question is missing something!!

OpenStudy (anonymous):

How many pizzas would need to be sold to make a profit of $1,100? How many pizzas would need to sell if you wanted to make a profit greater than $1,595

OpenStudy (hoblos):

so replace P=1100 and calculate x for the other one 11x -55 >1595 also solve for x can you do them by yourself??

OpenStudy (anonymous):

I will try and thanks for the help

OpenStudy (hoblos):

0k.. if you face any problem tell me and i'll try to help ;)

OpenStudy (anonymous):

I'm having problems solving this.